Please help me find an equation that is perpendicular to y=-5x+7 and passes through (10,2). Thank you!
Answer by stanbon(75887)

Please help me find an equation that is perpendicular to y=-5x+7 and passes through (10,2).
------
Given line has slope = -5
-----
Perpendicular line must have slope = 1/5
-------
Form:: y = mx + b
2 = (1/5)10 + b
2 = 2 + b
b = 0
-----
Equation:
y = (1/5)x
===================
Cheers,
Stan H.
================